Sir, can u help me with this one....
Near the surface of the earth, there exists a downward directed electric field of 150V/m. Using the Gauss theorem, calculate the surface charge density of the earth's surface. Assume earth to be a conducting medium.

At a height of 200m above the surface of earth, the downward field is 100V/m. Calculate the average volume charge density in the earths atmosphere between the earth's surface and 200m above the surface.

Express your answer as an excess or deficit of electrons per unit area or volume.....

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

E·dA=Q/k,where k=ε0.

because of symmetry, 4πr2*150=Q/k, E is downward.Thus, Q has negative sign.

surface charge density=Q/(4(pi)r2)=-150k